1) The approximate cell yield of a T25 flask (for adherent cells) is 2.8 x 106 cells. The cell yield is the total amount of adherent cells at confluence. How many adherent cells would you expect to be in this flask at 80% confluence?

2) For adherent cells, the seeding density is usually between what range of % confluencies?

3) How many total cells would you place in a T75 flask with a cell yield of 8.4 x 106 cells to seed the cells at 20% confluence for an experiment?

4) After observing your adherent cells in the T75 flask under the microscope, you estimate the confluence of the cells to be around 70%. When you passage the cells, you count the cells using a hemocytometer and obtain a total cell count of 6.5 x 106 cells. (T75 cell yield = 8.4 x 106 cells)

  1. What does the cell count (from the hemocytometer) suggest is the % confluence of the flask?
  2. Which number would you trust to be more accurate? Why?
  3. When doing experiments, which method of determining confluency/cell counts would result in the proper seeding densities in well-plates for an experiment

5) You have cells at 80% confluence in a T75 flask and you want to seed as many 12-well plates as possib

6) Continuing from the previous question . . .

  1. How would you resuspend the cells from the T75 after centrifugation to obtain a resuspension concentration of 2 x 105 cells per 0.1 mL?
  2. Each well in a 12-well plate will generally hold 800 µL of media. What volume of cells from your resuspension and what volume of media would you add to each well?
  3. What type of pipette would you use to transfer the media and cells to the well
Answer #1

1. 100% confluency = 2.8 * 10^6 cells

80% confluency = {80 * (2.8 * 10^6)} / 100 = 2.24 * 10^6 cells

2. Seeding density for adherent cells in a T25 flask = 0.7 * 10^6 cells

In terms of confluency it is = {(0.7 * 10^6) * 100} / (2.8 * 10^6) = 25 % confluency